### v3

# Visual Whole-Body Control for Legged Loco-Manipulation

Abstract
We study the problem of mobile manipulation using legged robots equipped with an arm, namely legged loco-manipulation. The robot legs, while usually utilized for mobility, offer an opportunity to amplify the manipulation capabilities by conducting whole-body control. That is, the robot can control the legs and the arm at the same time to extend its workspace. We propose a framework that can conduct the whole-body control autonomously with visual observations. Our approach, namely Visual Whole-Body Control (VBC), is composed of a low-level policy using all degrees of freedom to track the end-effector manipulator position and a high-level policy proposing the end-effector position based on visual inputs. We train both levels of policies in simulation and perform Sim2Real transfer for real robot deployment. We perform extensive experiments and show significant improvements over baselines in picking up diverse objects in different configurations (heights, locations, orientations) and environments.
Figure 1: Real-Robot input visualization and grasping trajectory. Our framework enables the robot to grasp different objects in varying heights and surroundings. Our policy is trained in simulation using visual inputs (masks and masked depth images from two views) and can be deployed in real robots without any real-world fine-tuning.
00footnotetext: The first two authors contributed equally.
I Introduction
The study of mobile manipulation has achieved large advancements with the progress of better manipulation controllers and navigation systems. While installing wheels for a manipulator can help solve most household tasks [75, 23], it is very challenging to adopt these robots outdoors with challenging terrains. Imagine going camping; having a robot picking up trash and wood for us would be very helpful. To achieve such flexibility and applications in the wild, we study mobile manipulation with legged robots equipped with an arm, i.e., legged loco-manipulation. Specifically, we are interested in how the robot can conduct tasks based on its visual observation autonomously.
An important aspect of loco-manipulation is the opportunity to conduct whole-body control using all degrees of freedom of the robot. While the legs are typically utilized for mobility, there is a large potential for controlling all the joints of the legs and arms together to amplify manipulation capabilities. We humans do this, too: We use our legs to stand up to reach a book placed on a shelf; We bend our legs to tie our shoes. Fig. 1 shows the tasks our robot can do when controlling the whole-body joints together. For example, our robot can go to the side of the road and pick up the trash on the grass. This will not be achievable without bending the front legs of the robot. The arm alone will not be long enough to reach the ground.
While this is appealing, it is a very challenging control problem to coordinate all the joints simultaneously with large degrees of freedom (19 DoF). The robot will need to exploit the contact with its surroundings and the objects and maintain stability and robustness to external disturbances all at the same time. Besides, acquiring a mobile robot platform to do precise manipulation jobs requires a steady robot body, which is much harder for legged robots compared to wheeled robots. An even larger challenge comes from achieving all these in diverse environments autonomously, given only the observation from an egocentric camera. Recent learning-based approaches have shown promising results on legged robots avoiding obstacles, climbing stairs, and jumping over stages robustly with visual inputs [13, 1, 82, 15]. However, all these efforts still focus only on locomotion without manipulation, which requires more precise control. The extra complexity and more challenging tasks make direct end-to-end learning infeasible.
In this paper, we conduct loco-manipulation by introducing a two-level framework with a high-level policy proposing end-effector pose and robot body velocity commands based on visual observations and a low-level policy tracking these commands. Such a hierarchical design is effective as the low-level controller can be effectively shared when interacting with diverse environments and objects. We named our framework Visual Whole-Body Control (VBC). Specifically, VBC is a hierarchical solution that contains three stages of training: first, we train a universal low-level policy using reinforcement learning (RL) to track any given goals and achieve whole-body behaviors; then, we train a privileged teacher policy by RL to provide a set of appropriate goals and guide the low-level policy to accomplish specific tasks (e.g., pick-up); finally, to deploy the policy into the real robot, we distill the teacher policy into a depth-image-based visuomotor student policy via online imitation learning. All this training is done in simulation, and we perform direct Sim2Real transfer for deployment in the real robot.
Our hardware platform is built on a Unitree B1 quadruped robot equipped with a Unitree Z1 robotic arm. We design a server-client multi-processing system that consists of four different modules working under different frequencies. We conducted the pickup task with 14 different objects, including regular shapes, daily uses, and irregular objects, under three height configurations, i.e., on the ground, on a box, and on a table, achieving a high success rate in all scenarios.
We summarize the contribution of this paper as:
  • We develop an autonomous vision-based loco-manipulation system that utilizes model-free RL and sim-to-real techniques to learn whole-body control policies for legged robots.
  • Our proposed method emerges retrying behavior and generalizes well to complicated terrains, various objects, and heights.
II Related Work
Legged Locomotion. For decades, legged locomotion has been an essential topic in the field of robotics and has gained much success. Classical methods solve the legged locomotion task through model-based control to define controllers in advance. For example, Miura and Shimoyama [50], Sreenath et al. [61] utilized traditional methods to enable the walking of biped robots. MIT cheetah 3 [7], ANYmal-a [31] by ETH, and Atlas [19] by Boston Dynamics are all perfect examples of robots designed with a robust walking controller. Based on these developed robots, researchers have been able to develop advanced controllers for even more agile locomotion tasks, such as jumping or walking over varying obstacles [53, 51, 26, 52, 3, 38]. To make a robust controller that can adapt to varying environments, researchers integrate some sensor data such as elevation maps [18, 40, 35, 6, 17] and state estimators [8, 66, 10, 74, 65]. Although classical methods have greatly succeeded in legged locomotion, generalizability is always a problem for pre-designed controllers, and acquiring these controllers requires careful and tedious engineering. With the fast development of deep neural networks, learning-based end-to-end methods have become a simple way to gain robust controllers with strong generalization abilities. Learning-based methods often train a robust controller in physics simulators such as PyBullet [14] and Isaac Gym [47] and then perform sim-to-real transfer to a real robot. For example, Rudin et al. [58] uses Isaac Gym with GPU to enable a quadruped robot to walk in only several minutes, Kumar et al. [36, 37] proposed the RMA algorithm to make the sim-to-real transfer process easier, and more robust, Lee et al. [41], Miki et al. [49], Agarwal et al. [1], Yang et al. [73], Yu et al. [80], Cheng et al. [13], Zhuang et al. [82], Fu et al. [21], Margolis and Agrawal [48], Li et al. [44] enable the legged robots to walk or run on varying terrains, and Fuchioka et al. [24], Escontrela et al. [16], Peng et al. [54], Yang et al. [71], Wang et al. [64], Li et al. [43], Kumar et al. [39] combine RL with imitation learning to produce naturalistic and robust controllers for legged robots. Kim et al. [33] map human motions to a quadruped robot to complete a wide range of locomotion and manipulation tasks.
Mobile Manipulation. Previously mentioned works mainly focus only on the mobility part; adding manipulation to the mobile base is yet to be well-solved. Autonomous mobile manipulation on wheeled robot platforms has made significant progress over the past few years. Wong et al. [68] collected teleoperated mobile manipulation data and learned imitation-learning-based policies in simulation. Lew et al. [42] built an efficient framework for table cleaning with RL and trajectory optimization. Garrett et al. [25] and Srivastava et al. [62] integrate task-specific planner for various mobile manipulation tasks. Gu et al. [27] and Xia et al. [69] developed modularized mobile manipulators with RL, Yokoyama et al. [78] introduced a high-level controller for coordinating different low-level skills. Sun et al. [63] designed a modularized policy with components for manipulation and navigation, trained by RL to learn navigation and grasping indoors. Yang et al. [72] proposed simultaneously controlling the robot arm and base for harmonic mobile manipulation using RL training with large-scale environments. Ahn et al. [2] pre-trained a set of single mobile or manipulation tasks and relied on large language models to decompose and call the skills to achieve long-horizon tasks. Brohan et al. [9] train a robotics transformer to perform real-world tasks given human instructions, trained on large-scale and real-world demonstration. Shafiullah et al. [59] learned to solve household tasks from 5-minute demonstrations based on pre-trained home representations. While encouraging, most of these mobile manipulators with wheels cannot operate with challenging terrains in the wild.
Legged Loco-Manipulation A set of works has exploited legged loco-manipulation to take advantage of legged robots. Bellicoso et al. [5] designs an online motion planning framework, together with a whole-body controller based on a hierarchical optimization algorithm to achieve loco-manipulation, Fu et al. [22] proposed an effective method to train a whole-body controller end-to-end through DRL. Ma et al. [46] combined model-based manipulator control and learning-based locomotion, but only showed the response to external push instead of achieving manipulation tasks. Notably, these works all require human involvement and teleoperation to execute and achieve tasks, rather than operating autonomously. Zhang et al. [81], Yokoyama et al. [79], Arcari et al. [4] all performed an autonomous mobile manipulation system on quadruped robots while they utilized the default model-based controller provided through the robot APIs (only control robot locomotion in a static height) and only focused on the manipulation part without any whole-body behaviors to adapt to different heights of objects. Ferrolho et al. [20] proposed a trajectory optimization framework based on a robustness metric for solving complex loco-manipulation tasks, but the task is assigned and the target positions are calculated via a motion capture system, thus only works limitedly. Sleiman et al. [60] realized a bilevel search strategy for motion planning, by incorporating domain-specific rules, combing trajectory optimization, informed graph search, and sampling-based planning, enabling robots to perform complex tasks. They require the user to provide a dense description of the scene to initialize the solver. Zimmermann et al. [83] also utilized trajectory optimization to enable a legged robot to walk and pick up objects. However, they only show limited surroundings and simple objects. Besides these works with arm-based configuration, Wolfslag et al. [67] use the legs for pushing and box lifting tasks by introducing additional support legs named prongs to enhance the robustness and manipulation capabilities of quadrupedal robots. Jeon et al. [32] adopted a similar hierarchical framework to us, but does not involve any visual inputs, and only learned a quadruped for pushing large objects without an additional arm.
The system we design highly differs from the aforementioned works, as our work is a fully autonomous and vision-based mobile manipulation system on quadruped robots, with limited human intervention. The robot is controlled and planned by trained policies and can adapt to objects with varying heights through its whole-body behavior, and emergent retrying behaviors when the task fails. Due to our design of visual inputs, our system can work both indoors and outdoors with no external limits, as illustrated in Fig. 1, compared with almost all of the above-mentioned works that are only tested in limited lab scenarios. In Fig. 2 we show why VBC is highly advantaged of the non-whole-body methods [81, 79, 4], for the body control behavior allowing much more flexible poses (e.g., bending) to reach objects in random heights, like the one on the ground.
(a) Non-whole-body method.
(b) VBC (ours).
Figure 2: Illustration of comparing VBC to non-whole-body method. The robot body of the non-whole-body method does not bend to allow the arm to reach the objects on the ground.
Figure 3: Training a vision-based whole-body manipulator. Our VBC framework trains a low-level control policy and high-level planning policy end-to-end separately for whole-body manipulation.
III Visual Whole-Body Control
In this section, we introduce our Visual Whole-Body Control (VBC) framework (as described in Fig. 3). Our VBC framework consists of a low-level goal-reaching policy and a high-level task-planning policy. Our low-level goal-reaching policy tracks a root velocity command and a target end-effector pose. Our vision-based high-level task-planning policy provides velocity and end-effector pose command to the low-level policy, given the segmented depth images and proprioception of the robot as inputs. When combined, our robot could operate fully autonomously using only visual input and its proprioception. We start with an overview of our robot setup, followed by the training pipeline of the low-level policy and the high-level policy, and our robot system design.
III-A Robot Platform Overview
The robot platform used in this work comprises a Unitree B1 quadruped robot with a Unitree Z1 robot arm and a gripper mounted on its back. Our robot is equipped with two cameras, one on the head and another mounted close to the gripper. The Unitree B1 quadruped has 12 actuatable DoFs, the arm has 6 DoFs, and the gripper has another 1 DoF, so there are 19 DoFs in total. An overview of the hardware setup is shown in Fig. 4.
Figure 4: Real robot system setup.
III-B Low-Level Policy for Whole-Body Goal-Reaching
The low-level control policy that takes over the whole-body locomotion control $\pi^{\text{low}}$ as described in the blue part in Fig. 3, is trained to track any given end-effector poses and body velocities across various terrains. In this work, we turn to the help of RL and utilize ROA (Regularized Online Adaptation) [22, 12, 32] to realize such a universal and robust behavior that can become the foundation of any high-level tasks.
Commands. The command $\mathbf{b}_{t}$ for our low-level policy is defined as:
$\mathbf{b}_{t}=[\mathbf{p}^{\text{cmd}},\mathbf{o}^{\text{cmd}},v_{\text{lin}}^{\text{cmd}},\omega_{\text{yaw}}^{\text{cmd}}]$
[…]
Figure 5: Comparison between trajectories and segmented depth images in simulation and real-world sampled by VBC. Left: The real-world behaviors match great with the simulation, indicating a less sim-to-real gap in our method. Right: Our TrackingSAM-based working pipeline allows us to obtain the segmentation mask precisely during deployment (not the same object observed). The depth images are pre-processed by hole-filling and clipping.
Observations. The observation of our policy is defined as a 90-dimensional vector $\mathbf{o}_{t}$:
$\mathbf{o}_{t}=[\mathbf{s}_{t}^{\text{base}},\mathbf{s}_{t}^{\text{arm}},\mathbf{s}_{t}^{\text{leg}},\mathbf{a}_{t-1},\mathbf{z}_{t},\mathbf{t}_{t},\mathbf{b}_{t}].$
Among them, $\mathbf{s}_{t}^{\text{base}}\in\mathbb{R}^{5}$ is the current quadruped base state including (row, pitch, and quadruped base angular velocities), $\mathbf{s}_{t}^{\text{arm}}\in\mathbb{R}^{12}$ is arm state (position and velocity of each arm joint except the end-effector), $\mathbf{s}_{t}^{\text{leg}}\in\mathbb{R}^{28}$ is leg state (joint position and velocity of each leg joint, and foot contact patterns that are 00 or $1$), $\mathbf{a}_{t-1}\in\mathbb{R}^{12}$ is the last output of policy network, $\mathbf{z}_{t}\in\mathbb{R}^{20}$ is environment extrinsic vector, which is an encoding of environment physics. Besides, following the work of [1], to learn a steady walking behavior (like trotting), we provide some extra timing reference variables $\mathbf{t}_{t}$ to the observation, which is computed from the offset timings of each foot by repeat the gait cycle of the desired symmetric quadrupedal contact pattern.
Actions. Our low-level goal-reaching policy controls the quadruped robot and the mounted arm simultaneously. For the quadruped robot, our low-level policy outputs the target joint angles for all 12 joints of the robot, which is converted to torque command by a PD controller.
Arm control. Previous work [22] has revealed that whole-body RL policy is good at controlling the end-effector position but struggles in controlling the end-effector orientation. Unfortunately, such behaviors are critical for our manipulation tasks. Thereafter, we take an off-the-self Inverse Kinematic (IK) Solver to convert the end-effector pose command into target joint angles for controlling the arm. Formally, IK computes the increment of the current joint position $\Delta\mathbf{\theta}\in\mathbb{R}^{6}$
$\Delta\mathbf{\theta}=J^{T}\left(JJ^{T}\right)^{-1}\mathbf{e},$ (1)
where $J$ is the Jacobian matrix, and $\mathbf{e}\in\mathbb{R}^{6}$ is the difference between the current end-effector pose and the next end-effector pose represented as the position and the Euler angle.
Reward function. Our reward for training the low-level policy for whole-body control mainly consists of four parts: command-following reward, energy reward, alive reward, and phase reward. The command following reward encourages the robot to track and follow the provided commands and explore various whole-body behaviors; the energy reward penalizes energy consumption to enable smooth and reasonable motion; the alive reward encourages the robot not to fail; the phase reward encourages the robot to walk in a steady behavior as mentioned above.
End-effector goal poses sampling. To train a robust low-level controller for goal-reaching, we sample and track various gripper poses. For sampling smooth trajectories of gripper target poses for tracking, we employ a height-invariant coordinate system originating at the robot’s base and randomly sample trajectories on this at every fixed timestep. This ensures the sampled trajectories are always on a smooth sphere and prevents the sampled goals from being affected by the quadruped robot’s height. Details can be found in Appendix II-A.
Policy architecture. The low-level control policy network is a three-layer MLP, where each layer has $128$ dimensions and ELU activations. It receives observations, including the extrinsic latent environment, which is encoded by the privileged encoder during training, and the adaptation module during testing. The privileged information encoder that models environment physics parameters is a two-layer MLP with a hidden size of $64$; the adaptation module encodes history rollouts, which is a two-layer convolutional network.
Domain randomization. When training, we randomize the type of terrains, including flat plane and rough ones. We also randomized the friction between robots and terrains. Besides, we randomized the quadruped’s mass and center of mass. These randomizations, together with our ROA module, contribute to a robust low-level policy.
III-C High-Level Policy for Task Planning
When deploying our policy into the real world, the robot can only perceive the object through depth images to achieve high-frequency control. However, RL with visual observation is quite challenging and requires particular techniques to be efficient [30, 77, 29, 28, 70]. To achieve simple and stable training, we choose to train a privileged state-based policy that can access the shape information of the object (as described in the orange part of Fig. 3) and distill a visuomotor student (as described in the green part of Fig. 3).
C.1. Privileged Teacher Policy
Like the low-level policy, the privileged teacher policy is trained through RL (we also use PPO in this work) with designed task-specific reward functions and domain randomization, with a fixed low-level policy underlying.
Privileged observations. The privileged observations include the encoded shape feature and object poses. We formally define it as a $1094$-dim vector:
$\mathbf{o}_{t}=[\mathbf{z}^{\text{shape}},\mathbf{s}_{t}^{\text{obj}},\mathbf{s}_{t}^{\text{proprio}},\mathbf{v}_{t}^{\text{base vel}},\mathbf{a}_{t-1}],$
where $\mathbf{z}^{\text{shape}}\in\mathbb{R}^{1024}$ is the latent shape feature vector encoded from the object point clouds using a pre-trained PointNet++ [55]. The pre-trained PointNet++ is fixed, so $\mathbf{z}^{\text{shape}}$ remains invariant during training. $\mathbf{s}_{t}^{\text{obj}}\in\mathbb{R}^{6}$ is the object pose in local observation w.r.t the robot arm base. It contains the local pose of the object. The use of local coordination inherently incorporates the information regarding the distance to the object. $\mathbf{s}_{t}^{\text{proprio}}\in\mathbb{R}^{53}$ consists of joint positions including gripper joint position $\mathbf{q}_{t}\in\mathbb{R}^{19}$, joint velocity not including gripper joint velocity $\dot{\mathbf{q}}_{t}\in\mathbb{R}^{18}$, and end-effector position and orientation $\mathbf{s}^{\text{ee}}_{t}\in\mathbb{R}^{6}$. $\mathbf{v}_{t}^{\text{base vel}}\in\mathbb{R}^{3}$ is the base velocity of robot and $\mathbf{a}_{t-1}\in\mathbb{R}^{9}$ is the last high-level action.
High-level Actions. The high-level policy $\pi^{\text{high}}$ outputs actions that determine the desired tracking position for the low-level policy. The high-level action is formally defined as:
$\mathbf{a}_{t}=[\mathbf{p}_{t}^{\text{cmd}},\mathbf{v}_{t}^{\text{cmd}},p_{t}^{\text{gripper}}]\in\mathbb{R}^{9},$
where $\mathbf{p}_{t}^{\text{cmd}}\in\mathbb{R}^{6}$ is the increment of gripper pose, $\mathbf{v}_{t}^{\text{cmd}}\in\mathbb{R}^{2}$ is the command of quadruped linear velocity and yaw velocity and $p_{t}^{\text{gripper}}\in\{0,1\}$ indicates whether the gripper should be close or open.
Figure 6: The server-client multi-processing system design. Modules work asynchronously under different frequencies. The high-level planning policy (left) and the low-level control modules (right, controlling arm and legs) are deployed on different devices. To be more specific, the goal receiver requests the high-level policy to take the visual input and receive the new commands; the low-level policy sends joint position commands to the quadruped runs at 50Hz, given the input of high-level commands; the IK-based arm controller runs and send the joint position command to the arm constantly at 800Hz, given the target end-effector pose. Therefore, the low-level policy and the arm controller keep utilizing the old commands to control the robot until the high-level policy updates the commands (i.e., target robot velocities and end-effector pose).
Reward functions. The reward function for training our state-based policy comprises stage rewards and assistant rewards, as defined in Tab. III. As for the task-specific rewards, we design three stages. The first stage is approaching, and the corresponding reward $r_{\text{approach}}$ encourages the gripper and quadruped to get close to the target object:
$r_{\text{approach}}=\min(d_{\text{closest}}-d,0)~{},$ (2)
where $d_{\text{closest}}$ is the current minimum distance between the gripper and the target object. The second stage is task progress, and the reward $r_{\text{progress}}$ leads the robot to the desired final state, e.g., lifting the object in the pickup task:
$r_{\text{progress}}^{\text{pickup}}=\min(d-d_{\text{highest}},0)~{},$ (3)
where $d_{\text{highest}}$ is the current largest height of the object. The third stage is denoted task completion, and reward $r_{\text{completion}}$ bonus when the agent completes the task.
$r_{\text{completion}}=\mathbf{1}(\text{Task is completed})~{},$ (4)
For example, in the pickup task, the task completion condition is that the object height exceeds the required threshold. Note that at one time, the robot can only be at one stage and receive one stage reward of the above three, ensuring that different task stages are not intervened in. The assistant rewards are designed to a) smooth the robot’s behavior to be easier to transfer to the real world, like penalty huge action commands and joint acceleration; b) prevent deviation and sampling useless data, such as acquiring the robot’s body and gripper to point to the object, which guarantees that the camera does not lose tracking objects. Beyond these major rewards, we also design a set of assistant reward functions that help accelerate the learning procedure, see details in Appendix II-B.
Domain randomization. When training our high-level policy, we randomized the friction between the robot and the terrain, the mass of the robot, and the center of mass of the robot. Although our low-level control policy can handle these gaps in the real world, its performances differ under different environmental conditions. These randomizations help the high-level policy adjust better to different low-level performances. Moreover, due to the uncertainty of the high-low signal transmission in the real world, we also randomize the high-level frequency by involving different low-level policy calls during one high-level call. We also randomize the Kp and Kd parameters for the arm controller as we use the built-in position control mode for the arm in the simulation. Regarding the pick-up task, we randomize table heights, the initial position and pose of the robot, along with the position and pose of objects.
C.2. Visuomotor Student Policy
We find robust and adaptive picking-up behaviors emerge from the privileged state policy training. Nonetheless, it is assumed that object poses and shape features, which are inaccessible during deployment, are taken as input. To deploy our policy in the real world, we need to use easily accessible real-world observations to infer (representations of) these properties.
Observations and actions. As camera images are easily accessible both in simulation and the real world, we use camera images that implicitly contain the information of object state and shape to train our teacher policy. To mitigate the deployment gap, we use only depth images, as RGB images in the real world tend to differ greatly from those in the simulator. Besides, we use segmented depth images instead of whole depth images to avoid the negative influence of surroundings. Formally, the observation of our vision policy is
$\mathbf{o}_{t}=[\mathbf{o}^{\text{image}}_{t},\mathbf{s}_{t}^{\text{proprio}},\mathbf{a}_{t-1}],$
where $\mathbf{o}^{\text{image}}_{t}$ consists of object segmentation masks and segmented depth images. $\mathbf{s}_{\text{proprio}}$ and $\mathbf{a}_{t-1}]$ are proprioceptive observation and last action defined above. The output actions are the same as those of state-based policy.
Imitation learning for student distillation. To deploy our policy in the real world, we distill our state-based high-level planning policy into a vision-based policy using DAgger. We warm up by teacher sampling the initial transition data. This is like behavior cloning (BC) to find a good initialization of the visuomotor student policy to improve the training efficiency. Then we sample transitions using the student policy and request correction from the teacher. We adopt augmentations like RandomErasing, GaussianBlur, GaussianNoise, and RandomRotation for the depth images. More details can be found in Appendix II-B.
Policy network architecture. Our vision-based high-level policy network comprises a CNN, a GRU, and a two-layer MLP that outputs the high-level actions. We concatenate the images and other states as observations. The images are passed through the two-layer CNN network with kernel size being $5$ and $3$, respectively, and then encoded as a $64$-dim latent vector. Then, we concatenate this latent vector with other states as a flattened vector. This vector is then passed to a one-layer GRU with hidden size 128, followed by a two-layer MLP with hidden size 64. We use ELU as the activation.
III-D Real World Deployment
Figure 7: Segmented examples of the TrackingSAM tool. At the beginning of each reset trial, we annotated the objects from the two RGBD cameras mounted on the gripper and the robot head. Then, TrackingSAM will keep tracking the segmented object and providing the mask.
Hardware setup. As visualized in Fig. 4, we mount two cameras on our robot platform, which are both RealSense D435 cameras. B1 itself has an onboard computer, and we also attached an additional onboard computer, Nvidia Jetson Orin, to the robot. The B1, Z1, and the Orin computer are all powered by B1’s onboard battery. The inference of the low-level policy and the control of the arm is conducted at the onboard computer of B1, and that of the high-level vision-based planning policy is made on Nvidia Jetson Orin. It is worth noting that the computation and the power of our robot system are fully onboard and, therefore, untethered.
System design. During deployment, the low-level policy that controls the quadruped works at a frequency of 50Hz; in the meantime, the high-level policy runs at 8Hz, which is five times slower than the low-level policy; besides, the arm is required to work at a frequency of more than 800Hz. Since these policies work on different computers, along with the quadruped and the arm should work under different frequencies, we design a client-server multi-processing system for this work, as shown in Fig. 6. We set up three processes at the low-level policy side: a quadruped robot control process, an arm control process, and a goal receiver process behaving as a client. The high-level policy combines the vision inputs from the depth cameras and the information from the low-level request as the observation works as a server. Every 0.1 seconds, the goal receiver sends a request with the necessary information to the server; the high-level policy makes the latest inference and sends the high-level action back to the robot; at the same time, the quadruped and the arm control processes gather the latest high-level commands to execute, while keep updating the necessary information for the goal receiver process.
Sim-to-real transfer. We made several efforts to bridge the sim-and-real gap. Specifically, we slightly randomize the position and rotation of both our cameras to enable our policy to handle slight camera offsets from simulation. Besides, we clip the depth image to a minimum of 0.2m to avoid low-quality images when the distance is too close to a real camera, after which we normalize the depth value. As the surrounding environment in the simulation and the real world are very different, to reduce the noise, we only use segmented depth images to train our visuomotor policy.
Vision masks. Note that our system works with only limited human intervention, i.e., we need to annotate and segment objects from environment backgrounds. To this end, we develop TrackingSAM [56], a tool that combines the tracking model AOT (Associating Objects with Transformers) [76] and SAM (Segment Anything) [34]. This allows us to annotate the object on the image from both cameras or an arbitrary image before starting the task and track that object to provide its real-time segmented masks. The tool is integrated into the high-level server, and a demo example can be found in Fig. 7. During real-world experiments, we annotated the objects from the two RGBD cameras mounted on the gripper and the robot head at the beginning of each reset trial. TrackingSAM will keep tracking the segmented object and providing the mask. The depth images obtained from the RGBD cameras are pre-processed by hole-filling and clipping. In real-world experiments, as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to both views to initialize an automated pickup.
IV Experiments
We conduct a set of experiments on pick-up tasks to compare VBC with baselines both in simulation and the real world, showing effectiveness from the following perspectives:
  • VBC solves the mobile manipulation (pickup) task over varying objects with different heights;
  • VBC has the advantage over easy baseline methods, such as decoupling mobile manipulation problems into separate moving and arm-control;
  • The object shape feature produced by pre-trained PointNet++ boosts learned policy grasp varying objects;
Experiment setups.
Figure 8: Illustration of simulation / real-world setup.
The simulation / real-world setup is shown in Fig. 8. Particularly, in the simulation, the high-level policy is trained to pick up various objects on a range of table heights $[0,0.65]$ meter. We utilize IsaacGym [47] for parallel training in a large set of environments. In the real world, we initialize objects in a random location and orientation, and put them in front of the robot where it is visible through both cameras (around 0.5$\sim$1.5m). We test three different heights by putting the objects on the ground, on a box, and on a table, corresponding to the heights of 0.0m, around  0.3m,  0.5m.
Figure 9: Examples of training objects and real-world objects.
Figure 10: Success rates of different student policies at different heights, tested in the simulator. Each distinct object for more than 100 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps ($\sim$ 18 seconds). The dots represent the mean performance of all objects, the upper and the lower bound represent the maximum and the minimum success rate overall 7 categories, where we also note their category name.
IV-A Simulation Results
Objects. The objects used for the benchmark problems are shown in Fig. 9. Specifically, we included $34$ objects in total, mainly adapted from the YCB dataset [11]. According to the object shapes, we roughly divided them into $7$ categories: ball, long box, square box, bottle, cup, bowl and drill.
TABLE I: Success rates of VBC compared to baseline method on seven categories of objects, tested in the simulator. Each distinct object for more than 300 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps. With the 3D shape feature, the algorithm is better at handling irregular objects. Detailed descriptions of each method can be referred to Section IV-A.
Policy Type Ball Long Box Square Box Bottle Cup Bowl Drill
Privleged Teacher Floating Base $63.25\%$ $20.51\%$ $72.66\%$ $38.07\%$ $69.55\%$ $71.88\%$ $4.17\%$
VBC w.o. Shape Feature $\mathbf{96.18}\%$ $74.26\%$ $\mathbf{96.80}\%$ $\mathbf{82.46}\%$ $84.38\%$ $55.84\%$ $73.77\%$
VBC (Ours) $89.08\%$ $\mathbf{92.83\%}$ $81.87\%$ $76.34\%$ $\mathbf{89.12\%}$ $\mathbf{78.14\%}$ $\mathbf{84.00\%}$
Visuomotor Student Non-Hierarchical $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$
Floating Base $42.45\%$ $0.0\%$ $37.36\%$ $8.33\%$ $41.41\%$ $43.57\%$ $3.70\%$
VBC (Ours) $\mathbf{55.40}\%$ $\mathbf{28.57}\%$ $\mathbf{80.00}\%$ $\mathbf{56.57}\%$ $\mathbf{68.01}\%$ $\mathbf{58.96\%}$ $\mathbf{53.33\%}$
(a) Objects on the ground (0.0m).
(b) Objects on a box ($\sim$0.3m).
(c) Objects on a table ($\sim$0.5m)
Figure 11: Average success times w.r.t. the trial times of picking up 14 objects at different heights, tested in the real world. We allow 5 continuing trial times (including 4 continuous retries, x-axis) during one experiment, i.e., one reset of picking up one specific object. A failure is recorded if the robot attempts more than 5 trails in one reset or if the object falls from a higher surface. For each object and each height, such experiments are conducted 5 times, and therefore we calculate the average success time (y-axis) across these trials, representing the mean number of success times across five trials. It is obvious that the emergent retrying behavior improves the performance.
Evaluation principles and baselines. In the simulation, when the environment is reset and a new trail begins, both the robot and the objects are initialized with random positions and poses, and it’s considered a success if an object is picked up and a failure when the object falls down the table or is not successful in 150 high-level steps. During experiments, we compare both the privileged teacher policies of VBC against baselines, and their visuomotor student policies over $34$ kinds of different objects: 1) VBC w.o. shape feature: a visuomotor policy that is trained the same as VBC, but does not access the pre-trained object shape features during the teacher training stage. This aims to show the necessity of the object shape features during teacher training. We only compare the teacher’s performance as the only difference comes from the teacher’s side. 2) Floating base: a floating base policy with a perfect low-level navigation ability but without whole-body behavior; in other words, the robot can always follow the given velocity commands and body height commands ranging from 0.4 to 0.55 meters (we set 0.4 meters as the lower bound as the default controller in real-world only allows 0.47 meters in minimum). This is to show the indispensability of the whole-body behavior. We compare both the teacher and its corresponding student policy using the same policy network structure. 3) Non-hierarchical: a unified policy trained with low-level policy and visuomotor high-level policy jointly in an end-to-end style. This policy takes the observation of our low-level and high-level policies together. It outputs all the target positions of 12 robot joint angles and the target pose of the gripper. We failed in training such a policy, demonstrating that the training pipeline of VBC is inevitable.
Figure 12: Visualization of robot behavior and corresponding commands during the test time. Particularly, we plot two kinds of high-level commands given by the high-level policy, as described in Section III-C, i.e., the velocity commands (including the quadruped linear velocity and yaw velocity); along with the gripper command (indicating whether the gripper should be close or open).
Picking up different objects. We test the picking-up success rate on each distinct object for more than $300$ episodes and collect the results over every category. During test time, all object heights are randomly set from $0.0$m to $0.6$m for each trial. As is shown in Tab. I, VBC achieves the best performance on 4/7 categories of objects. It is also interesting that with 3D features, VBC becomes worse at regular objects (compared to VBC w.o. Shape Feature), but is much better on irregular objects with complicated shapes, i.e., long box, cup, bowl and drill. One possible reason is that, without pre-trained features, the policy targets the object’s center for grasping, which is effective enough on simpler objects yet inapplicable for objects with more complicated shapes. One may notice that the student policy is much worse than the teacher policy, this is because when the robot touches but fails to grasp on the first try, the object will be randomly laid so that its pose will be much harder to grasp within only 150 steps using only visual inputs. In addition, the performance of the student policy of VBC behaves the worst on long boxes. This is because without shape features, it is much harder to determine the pose of the object, and some of the initial poses of long boxes are extremely hard to grasp for the gripper we used at some height (also as shown in Fig. 10).
Picking up over various heights. We test the mean success rate on each object category on a fixed height, where we also list the object with the maximum/minimum success rate on each height. We compare VBC and the floating base baseline on 7 different heights, and illustrate the results in Fig. 10. From the significant improvement of VBC on almost every height, compared to the floating base policy, we highlight the advantage of VBC that is able to achieve flexible whole-body behaviors in achieving high-level tasks.
IV-B Real-World Experiments
Objects In the real world, we choose 14 objects, including four irregular objects, six common objects, and four regular shapes, shown in Fig. 9. It is worth noting that almost all objects are unseen in the training, i.e., there are no identity objects during real-world experiments.
Evaluation principles and baselines. We deploy the trained visuomotor student policy of VBC as tested in the simulator directly into the real world. We In the real world, we test how many times the robot can pick up the assigned object in 5 trials (i.e., the robot is allowed to continuously retry four times when it fails to grasp without resetting), as we find that the robot emerges to have retrying behavior, as will be described below. We choose this as a measurement since we find that the emergent retrying behaviors improve the robustness, reliability, and success rate of finishing a job. We consider it a success when the object is picked more than 0.1m higher than the plane where it is placed and a failure when the robot tries more than five times or the object falls from its surface. Each object is evaluated over three different heights, as stated before, and five resets are permitted for each height. In this part, we compare VBC with a similar high-level policy trained with stationary legs. During the test of the real-world mobile manipulation task, we put the objects at an appropriate distance that the cameras mounted on the robot can observe, thus they can be manually annotated using our developed TrackingSAM tool. After the annotation, the robot is able to walk by and pick up the objects in a fully autonomous way.
Emergent retrying behaviors. It is worth noting that during simulation training, the robot emerges to learn a retrying behavior. In other words, when the robot fails to grasp an object, it will automatically retry to grasp without any human intervention. This shows an advantage and a clear difference compared to model-based methods.
Pickup performance. We collect all results and conclude the averaged performance over all objects on each height setting and show in Fig. 11. It is worth noting that our method allows for autonomous retrying behaviors when one pick-up trial fails. Obviously, VBC surpass the baseline method on all settings. To highlight, the baseline methods fail at 0.0m and 0.3m, as the default controller keeps a fixed robot height, similar to the floating base baseline. Even on the 0.5m setting, where the teleoperated baseline without whole-body behavior is equivalent to a static robotics arm, VBC is generally better. Besides, VBC shows great generalization ability on unseen shapes, thanks to the training strategy and observation design in the training pipeline.
Qualititive analysis. We want to analyze whether the commands sent by the high-level policy are reasonable to call the low-level policy to achieve the task. To this end, we visualize the velocity commands, the gripper command, along with their corresponding robot behavior during the testing time in Figure 12. It appears that there is less gap when deploying our policy trained in the simulator into the real world, which is able to provide the correct commands that show clear phases.
V Conclusions, Limitations and Future Works
In this paper, we proposed a fully autonomous mobile manipulation system based on quadruped robots and a hierarchical training pipeline, Visual Whole-Body Control (VBC). VBC consists of a low-level goal-reaching policy and a high-level task-planning policy, which is trained through reinforcement learning (RL) and imitation learning (IL). Although we showed great success in picking various objects, we found several limitations of the system: 1) Compounding error. Our pipeline is hierarchical and requires each module to be coherent and precise to make the robot work well. 2) In-precise depth estimation. The RGBD camera does not reflect a perfect depth estimation, especially on reflective objects. 3) Bad gripper. The gripper of the Unitree Z1 Arm is a beak-like gripper instead of a parallel one, which tends to push objects away and is hard for accurate manipulation. 4) Unstable vision inputs. The tracking SAM that is used to provide the mask of annotated objects, may lose track due to camera deviation and occlusion, or be confused when the selected objects have a similar color to the other things. In our current pipeline, before the robot autonomously picks up objects, we must manually specify the target object in both cameras mounted on the robot. Thanks to the generalizability of trackingSAM, we find that as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to another view for automated picking, and free the robot from being specified in the between of the range of the two cameras, which improves the usability of our system. Since the requirements of annotation and object visibility are totally decoupled from our sim-to-real framework, in the future we can further improve and replace these modules with automated annotation algorithms and a search/exploration module to further improve the automation of the whole system. Future works should also consider upgrading the hardware, like replacing a better gripper and RGBD cameras that provide better depth estimation; and simplify the working pipeline.
Acknowledgments
We would like to thank Unitree for their technical support. We thank Chengzhe Jia for the great help on 3D modeling. We also thank Haichuan Che, Jun Wang, Shiqi Yang, and Jialong Li for their kind help in recording demos.
Appendix
I Notations and Background
I-A Reinforcement Learning
Reinforcement learning (RL) is usually formulated as a $\gamma$-discounted infinite horizon Markov decision process (MDP) $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\rho_{0},r,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the environment dynamics function, $\rho_{0}:\mathcal{S}\rightarrow[0,1]$ is the initial state distribution, $r(s,a):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, and $\gamma\in[0,1]$ is the discount factor. The goal of RL is to train an agent with policy $\pi(a|s):\mathcal{S}\rightarrow\mathcal{A}$ to maximize the accumulated reward $R=\sum_{t}\gamma^{t}r(s_{t},a_{t})$.
I-B Proximal Policy Optimization
Proximal policy optimization (PPO) is one of the popular algorithms that solve RL problems. The basic idea behind PPO is to maximize a surrogate objective that constrains the size of the policy update. In particular, PPO optimizes the following objective:
$L^{PPO}(\theta_{\pi})=\mathbb{E}_{\pi}[\min(rA,\text{clip}(r,1-\epsilon,1+\epsilon)A)],$ (5)
where $r=\frac{\pi(a|s)}{\pi_{old}(a|s)}$ defines the probability ratio of the current policy and the old policy at the last optimization step, $A(s,a)$ is the advantage function, and $\theta_{\pi}$ is the parameter of the policy $\pi$.
I-C Imitation Learning
In general, imitation learning (IL) [57, 45] studies the task of learning from expert demonstrations. In this work, instead of learning from offline expert data, we refer to an online IL method, dataset aggregation (DAgger) [57] such that we request an online expert to provide the demonstrated action. Formally, the goal is to train a student policy $\hat{\pi}$ minimizing the action distance between the expert policy $\pi_{E}$ under its encountered states:
$\hat{\pi}=\arg\min_{\pi\in\Pi}\mathbb{E}_{s\sim d_{\pi}}[\ell(s,\pi)].$ (6)
In this paper, $\ell$ is the mean square error in practice.
I-D Whole-Body Control
Whole-body control in robotics refers to controlling a mobile manipulator with the function of locomotion and manipulation by one unified framework. In this work, we consider a system that contains a quadruped robot combined with a 6-DoF robot arm, each of which has an underlying PD controller, and we use a learned policy to generate control signals and feed them to the controller given 1) the robot states and 2) commands.
In detail, the policy used for controlling the quadruped robot is driven by external commands, such as the base velocity command; and the arm is controlled by a PD controller driven by the $\mathbb{S}\mathbb{E}(3)$ end-effector position-orientation command.
TABLE II: Details of low-level policy.
Definition of Symbols
Name Symbol
Leg joint positions $\mathbf{q}$
Leg joint velocities $\dot{\mathbf{q}}$
Leg joint accelerations $\ddot{\mathbf{q}}$
Target leg joint positions $\mathbf{q}^{*}$
Leg joint torques $\mathbf{\tau}$
Base linear velocity $v_{b}$
Base angular velocity $\omega_{b}$
Base linear velocity command $v_{x}^{*}$
Base angular velocity command $v_{\text{yaw}}^{*}$
Number of collisions $n_{c}$
Feet contact force $\mathbf{f}^{\text{foot}}$
Feet velocity in Z axis $\mathbf{v}^{\text{foot}}_{z}$
Feet air time $t_{\text{air}}$
Timing offsets $\mathbf{\theta}^{\text{cmd}}$
Stepping frequency $f^{\text{cmd}}$
Reward Function Details for the Low-Level Policy
Name Definition Weight
Linear velocity tracking $\phi(v_{b,xy}^{*}-v_{b,xy})$ $1$
Yaw velocity tracking $\phi(v_{\text{yaw}}^{*}-\omega_{b})$ $0.5$
Angular velocity penalty $-\|\omega_{b,xy}\|^{2}$ $0.05$
Joint torques $-\|\mathbf{\tau}\|^{2}$ $0.00002$
Action rate $-\|\mathbf{q}^{*}\|^{2}$ $0.25$
Collisions $-n_{collision}$ $0.001$
Feet air time $\sum_{i=0}^{4}(t_{air,i}-0.5)$ $2$
Default Joint Position Error $\exp{(-0.05\|\mathbf{q}-\mathbf{q}_{\text{default}}}\|)$ $1$
Linear Velocity z $\|v_{b,z}\|^{2}$ $-1.5$
Base Height $\|h_{b}-h_{b,\text{target}}\|$ $-5.0$
Swing Phase Tracking (Force) $\sum_{\text{foot}}\left(1-C_{\text{foot}}^{\text{cmd}}(\mathbf{\theta}^{\text{cmd}},t)\right)\left(1-\exp\left(-\|(\mathbf{f}^{\text{foot}})^{2}/\sigma_{cf}\|\right)\right)$ $-0.2$
Stance Phase Tracking (Velocity) $\sum_{\text{foot}}\left(C_{\text{foot}}^{\text{cmd}}(\mathbf{\theta}^{\text{cmd}},t)\right)\left(1-\exp\left(-\|(\mathbf{v}^{\text{foot}}_{z})^{2}/\sigma_{cv}\|\right)\right)$ $-0.2$
Simulation We use the GPU-based Isaac Gym simulator [47] where we train 4096 environments in parallel.
II Training Details
II-A Low-Level Training
Low-level commands sampling. To train a robust low-level controller for goal-reaching, we should sample and track various goals (i.e., base velocity commands and end-effector commands) as possible. As for the base velocity commands, we uniformly sample the linear velocity from a range of $[-0.6,0.6]$, and the angular velocity from a range of $[1.0,1.0]$. Regarding the end-effector, we employ a height-invariant coordinate system originating at the robot’s base. Such a design ensures the sampled trajectories are always on a sphere, guaranteeing the consistency of these samples; besides, it prevents the sampled goals from being affected by the quadruped robot’s height. In particular, we interpolate the end-effector goal command $\mathbf{p}^{\text{cmd}}$ between the current end-effector position $\mathbf{p}$ and a randomly selected end-effector position $\mathbf{p}^{\text{end}}$, at intervals of $T_{\text{traj}}$ seconds:
$\mathbf{p}_{t}^{\mathrm{cmd}}=\frac{t}{T_{\text{traj}}}\mathbf{p}+\left(1-\frac{t}{T_{\text{traj }}}\right)\mathbf{p}^{\text{end}},t\in\left[0,T_{\text{traj }}\right].$
$\mathbf{p}^{\text{end}}$ is resampled in a fixed timestep or if any $\mathbf{p}_{t}^{\text{cmd}}$ leads to self-collision or collision with the ground, similar to the linear and yaw velocity commands $v_{\text{lin}}^{\text{cmd}}$, $\omega_{yaw}^{\text{cmd}}$ that are uniformly sampled from a given range. Fig. 13 shows a screenshot of the low-level training, where the robot is required to track the given velocity commands and the end-effector pose.
Figure 13: A screenshot of low-level training, where the robot is required to track the given velocity commands (linear and yaw) and the end-effector pose.
Regularized online adaptation for sim-to-real transfer. We take the training paradigm of Regularized Online Adaptation (ROA) [22, 12, 32] in the RL training to help decrease the realizability gap when deploying our system into the real world. As illustrated in Fig. 3, we have an encoder $\mu$ that takes the privileged information $e$ such as friction and mass as input and predicts an environment extrinsic latent vector $z^{\mu}$. Then, the adaptation module $\phi$ estimates this $z^{\mu}$ only based on recent history low-level observations. The adaptation module $\phi$ is trained by imitating $z^{\mu}$ online. We can then formulate the loss function of our whole low-level policy w.r.t. policy’s parameters $\theta_{\pi^{\text{low}}}$, privileged information encoder’s parameters $\theta_{\mu}$, and adaptation module’s parameters $\theta_{\phi}$ as:
$\begin{split}L\left(\theta_{\pi^{\text{low}}},\theta_{\mu},\theta_{\phi}\right)=-L^{PPO}(\theta_{\pi^{\text{low}}},\theta_{\mu})&+\lambda\left\|z^{\mu}-\operatorname{sg}\left[z^{\phi}\right]\right\|_{2}\\ &+\left\|\operatorname{sg}\left[z^{\mu}\right]-z^{\phi}\right\|_{2}\end{split}$ (7)
where $\operatorname{sg}\left[\cdot\right]$ is the gradient stop operator, and $\lambda$ is the Laguagrian multiplier. We optimize this loss function using the dual gradient descent method, and [22] has identified that such an optimization converges under mild conditions. It is worth noticing that RMA [36] is a special case of ROA, where the Laguagrian multiplier $\lambda$ is set to be constant zero. The adaptation module $\phi$ starts training only after convergence of the policy $\pi^{\text{low}}$ and the encoder $\mu$.
Reward functions. The detailed definitions of the low-level reward functions are shown in Tab. II, where we define $\phi=\exp(-\frac{\|x\|^{2}}{0.25})$ for tracking the desired velocity of the robot. It is worth noting that we introduce the swing phase tracking rewards to help the robot learn trotting, and $C_{\text{foot}}^{\text{cmd}}(\theta^{\text{cmd}},t)$ computes the desired contact state of each foot from the phase and timing variable, as described in [48].
II-B High-Level Training
The privileged teacher policy is trained through RL, specifically, we use PPO for training with designed reward functions. Reward functions We put all reward functions along with their weights of training the high-level policies in Tab. III. Except for $r_{\text{progress}}$, $r_{\text{completion}}$ and $r_{\text{completion}}$ that is determined by the task stage, we define a set of assistant rewards to help the algorithm converge faster. Among them, $r_{\text{acc}}$ limits the arm joints velocities’ change rate; $r_{\text{action}}$ helps smooth the output of the high-level policy to prevent it from changing quickly; $r_{\text{ee orn}}$ and $r_{\text{base orn}}$ encourages the arm and the robot body to be aimed at the object; $r_{\text{base approach}}$ encourages the robot to be close to the object when the robot is far away from the object. Beyond these rewards, we also design curriculums such as gradually adding some rewards, commanding the low-level agent using the action of the last step considering the delay on the real robot, and clipping the high-level action outputs to help safely real-world deployment.
TABLE III: Details of High-Level Rewards, where $\dot{\mathbf{q}}$ denotes arm joint velocities; $v_{x}^{*}$ denotes base velocity command; $\mathbf{d}_{\text{obj}}$, $\mathbf{d}_{\text{ee}}$ and $\mathbf{d}_{\text{base}}$ are the direction from robot body to the object, the direction from end-effector to the object and the direction of the robot body orientation, separately; $\mathbf{x}_{\text{obj}}$ and $\mathbf{x}_{\text{base}}$ are the position coordinates in world frame of the object and the robot body.
Definition Weights
Stage Rewards $r_{\text{approach}}$ (Eq. (2)) $0.5$
$r_{\text{progress}}$ (Eq. (3)) $1.0$
$r_{\text{completion}}$ (Eq. (4)) $3.5$
Assistant Rewards $r_{\text{acc}}=1-\exp(\|\dot{\mathbf{q}}_{t-1}-\dot{\mathbf{q}}_{t}\|)$ $-0.001$
$r_{\text{cmd}}=-\|v_{x}^{*}+0.25\exp(-\|v_{x}^{*}\|)$ $1.0$
$r_{\text{action}}=1-\exp(-\|\mathbf{a}_{t-1}-\mathbf{a}_{t}\|)$ $-0.001$
$r_{\text{ee orn}}=|\mathbf{d}_{\text{obj}}\cdot\mathbf{d}_{\text{ee}}|$ $-0.05$
$r_{\text{base orn}}=\text{cos}(\mathbf{d}_{\text{obj}},\mathbf{d}_{\text{base}})|$ $0.25$
$r_{\text{base approach}}=|\mathbf{x}_{\text{obj}}-\mathbf{x}_{\text{base}}|$ $0.01$
Simulation We use the GPU-based Isaac Gym simulator [47]. As for the teacher, we train 10240 environments in parallel; regarding the student, we only utilize 240 parallel environments as there is a huge GPU memory cost for rendering visual observations. Experiments are trained on a single GPU (we use GTX4090 and GTX3090). On a GTX4090, The teacher policy takes around 36 hours for training, and the student policy consumes around 48 hours.
Additional Training Techniques We design several training techniques with our prior knowledge to help the high-level policy to learn a well-performed, stable, not only in the simulation, but can also be transferred into the real world. We briefly introduce them in the following:
  1. 1.
    Action delay. We add a one-step action delay to cover the inference and execution in the real world.
  2. 2.
    Command clip curriculum. We gradually clip the linear velocity command during training, as we find that a large velocity helps accelerate the learning convergence but a slow velocity highly stabilizes the real world performance.
  3. 3.
    Reward curriculum. The command penalty reward $r_{\text{cmd}}$ is designed to encourage the robot to slow its velocity when it is close to the object to be picked.
  4. 4.
    Randomly changes the object position. During training, we take a small probability ($10\%$) to randomly change the object position and pose. This helps the robot to learn the retrying behaviors when it finds it fails during one trial.
  5. 5.
    Forcing stop when closing gripper. We forcibly set the velocity command to 0 when the gripper command is closing. This contributes a lot to a stop-then-pick behavior and thus improves the performance.
These training techniques are adopted for both teacher policy and student policy training. When training the student policy, we additionally randomize the camera latency as the policy will never obtain the immediate images at the time it infers, but some pictures that are a few milliseconds before, similar to the real world. We randomize the value by measuring the camera latency of the device we used in our real-world experiments.